OpenStudy (anonymous):

Write the slope intercept equation for the line that passes through (-12,10) and is perpendicular to 4x+6y=4 Show all work please

OpenStudy (anonymous):

4x+6y=4 y=4/6 - 4x/6 y=2/3 - 2x/3 perpendicular lines m1m2=-1 -2/3 x m2 = -1 m2= 1.5 y-y1 = m(x-x1) y-10 = 1.5(x--12) y-10 = 1.5x+18 y=1.5x +28